# Operation of silicon photomultipliers in a dilution refrigerator down to 9.4 mK towards a cryogenic cosmic ray muon veto system

### v1

@@ 4 Towards a cryogenic cosmic ray muon veto system
The main motivation for testing the NUV-HD-cryo SiPM technology in an an ultra-low temperature environment is to determine whether it is feasible for SiPMs to be deployed as photon sensors for the cryogenic cosmic ray muon veto system currently being designed for the QUEST-DMC experiment. As outlined in Section 1, the concept design makes use of a scintillator volume surrounding the bolometer box holding the superfluid 3He target, coupled to a SiPM. High energy cosmic ray muons are minimum ionising particles which deposit approximately 2 MeV/g/cm2 of energy in materials they traverse. If a cosmic ray muon passes through QUEST-DMC, it will also deposit a fraction of its energy into the scintillator surrounding the experiment, thus producing a flux of scintillation photons. These can be used as a veto signal, to remove any fiducial volume events from the dark matter search that occur in coincidence with a scintillator signal consistent with a cosmic ray muon. If the SiPM is directly coupled to the scintillator (i.e., without using fibres to optically couple the physically separated scintillator and SiPM), it will need to be capable of operating at the same temperature as the scintillator, which is thermally coupled to a thermal radiation shield at a temperature of $\mathcal{O}$(10s mK).
@@
In addition to the main dataset used for the characterisation analysis, we also acquired a second, shorter dataset where a small volume of scintillator was coupled to the SiPM. We acquired 2500 waveforms per voltage under the same trigger conditions as before, with a shorter 10 µs acquisition window (1 µs pre-trigger, 9 µs post-trigger). The plastic scintillator used for this study was an early prototype for the SoLid experiment [abreu2018optimisation], and is reported to have a scintillation efficiency of approximately 10000 photons/MeV e- [Eljen]. This was a convenient test sample for the purpose of these initial proof-of-concept measurements.
@@
Three small blocks of the scintillator, each with dimensions of 1 cm (W) $\times$ 3 cm (L) $\times$ 0.5 cm (D), are attached to the inside of a spare copper box lid using Araldite. The scintillator blocks are each embedded with three wavelength shifting fibres [WLSfibres], which shift blue light into green light and guide the photons produced in the scintillator towards the face of the SiPM. Figure 13 shows a picture of the scintillator blocks attached to the inside of the copper box lid. Whilst the SiPM is not physically in contact with the scintillator blocks when the box is closed, there is an extremely small gap of 1-2 mm between the scintillator blocks and the SiPM face when the box is closed. The SiPM sits in the central area of the three scintillator blocks, for maximum coverage and light collection efficiency.
@@
Figure 13: Image showing the three scintillator blocks attached to the inside of the spare copper box lid. Each block is embedded with three optical fibres, which guide photons generated in the scintillator towards the SiPM surface.
@@
Figure 14 (left) shows an example event observed in this dataset, acquired at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V). In this event, the number of photons incident on the SiPM is so large that it surpasses the dynamic range of the oscilloscope, illustrated by the clipping of the signal at 140 mV. For high photon fluxes such as this, we also observe a lengthening of the signal tail, compared to the SiPM recharge time (<<1 µs). This is caused by the superposition of pulses from real photons, optical crosstalk, and afterpulsing, which contribute to an overall increase in the duration of the signal.
@@
Figure 14 (right) compares the pulse charge distribution, computed as the 10 µs integral about the event trigger, for data acquired with the scintillator (red) versus data acquired without the scintillator (blue). Without scintillator, we only expect to observe signals associated with the SiPM’s intrinsic noise, resulting in a narrow distribution peaked at the single-photon charge. However, once scintillator is introduced, we observe a much broader distribution and a high charge tail, corresponding to energy depositions in the scintillator from particle interactions. This distribution is likely a combination of environmental gamma-rays, which dominate the lower part of the charge distribution, and high-energy cosmic ray muons populating the high charge tail, as observed in the cryogenic cosmic ray muon veto system developed by the NUCLEUS experiment [NUCLEUS]. This demonstrates that the SiPM is clearly detecting light from the scintillator and optical fibres, which is easily distinguishable above the SiPM noise sources.
@@
As discussed in Section 3.4.1, the SiPM exhibits an increase in noise from long-lasting APs when operated at ultra-low temperatures. Whilst this additional noise source clearly limits the photon-counting capability of the SiPM for a low photon flux, this is less of a concern for high photon fluxes generated from cosmic ray muons. An increase in long-lasting APs will manifest as an increase to the integrated charge of a signal, but only in response to a primary energy deposition in the scintillator. As such, this should not limit the ability of the SiPM to tag cosmic ray muons through their initial energy deposition. This effect will however need to be fully characterised in order to model the observed charge distribution and determine an appropriate veto threshold. This is a first proof-of-concept of using scintillator directly coupled to a NUV-HD-cryo SiPM in a mK environment as a means of identifying cosmic ray muons.
@@
Figure 14: Left: Example waveform trace taken at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V), illustrating a candidate cosmic ray muon event. Right: Pulse charge distribution (integrated over a 10 µs window) about the event trigger at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 4.8 V) when SiPM is not coupled to scintillator (blue) versus when SiPM is coupled to scintillator (red).
@@ 5 Conclusions
We have shown that FBK NUV-HD-cryo SiPMs are capable of operating successfully in a dilution refrigerator, down to a base temperature of 9.4 $\pm$ 0.2 mK. We measure the reverse I-V curve of a single SiPM at room temperature and at the base temperature, $T_{\mathrm{MC}}$, determining the breakdown voltage to be $V_{\mathrm{bd,RT}}$ = 32.18 $\pm$ 0.07 V and $V_{\mathrm{bd,MC}}$ = 27.16 $\pm$ 0.05 V respectively. The observed decrease in the breakdown voltage of the SiPM is consistent with the expected relationship between the breakdown voltage and temperature for this technology. We observe that there is no difference between the measured breakdown voltage at $T_{\mathrm{MC}}$ and the average breakdown voltage at 77 K of 27.19 $\pm$ 0.05 V measured by the DarkSide-20k collaboration for this technology [QAQC].
[…]
We measure the average number of correlated delayed avalanches, dominated primarily by afterpulses, for three different integration windows following a primary avalanche. The average number of CDAs is measured to be $\bar{N}_{\mathrm{CDA}}$ = 0.170 $\pm$ 0.002 (1 µs window), 0.420 $\pm$ 0.003 (5 µs window), and 0.560 $\pm$ 0.004 (10 µs window) at $\Delta V$ = 4.8 V. $\bar{N}_{\mathrm{CDA}}$ increases by 102% (1 µs window), 146% (5 µs window), and 172% (10 µs window) between $\Delta V$ = 4.8 V and $\Delta V$ = 6.8 V. The afterpulsing probability is reported to be 12% at 77 K [NUV-HD-cryo], compared to our measurement of 43% at $\Delta V$ = 4.8 V (assuming a 10 µs integration window). A significant increase in afterpulsing has previously been reported for SiPMs operating at temperatures below 40 K [hanski2025performance, zhang2022scintillation], which results in self-sustaining cascades of afterpulses that can last for up to approximately 1 ms at low temperatures. We observe evidence consistent with an increase in long-lasting afterpulses in our data, the effect of which is enhanced for higher $\Delta V$.
@@
Finally, we present initial measurements of deploying a NUV-HD-cryo SiPM coupled to a scintillator with optical fibre assembly, operated down to 9.4 mK, as proof of concept for a cryogenic cosmic ray muon veto system to be operated internal to a dilution refrigerator. In this run, we observe a significantly larger charge spectrum with a high charge tail, consistent with energy depositions from environmental gamma-rays and high-energy cosmic ray muons that are not present in the data acquired without the scintillator. The work presented in this manuscript will be used to inform the final design and optimisation of the cryogenic cosmic ray muon veto system currently under development for the QUEST-DMC experiment.

### v2

@@ 4 Towards a cryogenic cosmic-ray muon veto system
The main motivation for testing the NUV-HD-cryo SiPM technology in an an ultra-low temperature environment is to determine whether it is feasible for SiPMs to be deployed as photon sensors for the cryogenic cosmic-ray muon veto system currently being designed for the QUEST-DMC experiment. As outlined in Section 1, the concept design makes use of a scintillator volume surrounding the bolometer box holding the superfluid 3He target, coupled to a SiPM. High energy cosmic-ray muons are minimum ionising particles which deposit approximately 2 MeV/g/cm2 of energy in materials they traverse. If a cosmic-ray muon passes through QUEST-DMC, it will also deposit a fraction of its energy into the scintillator surrounding the experiment, producing a flux of scintillation photons. These can be used as a veto signal, to remove any fiducial volume events from the dark matter search that occur in coincidence with a scintillator signal consistent with a cosmic-ray muon. If the SiPM is directly coupled to the scintillator (i.e., without using fibres to optically couple the physically separated scintillator and SiPM), it will need to be capable of operating at the same temperature as the scintillator, which is thermally coupled to a thermal radiation shield at a temperature of $\mathcal{O}$(10s mK).
The requirements of the cryogenic cosmic-ray muon veto system are set by both its primary function of tagging muon-induced events, and the constraints of operation near the main experimental volume at mK temperatures. The system must provide a sufficiently large response to muon energy deposits to allow a practical threshold to be defined in the presence of background interactions from environmental gammas, targeting a veto efficiency of $\geq$90% while keeping accidental vetoes - and hence experimental dead time - minimal. A detailed characterisation of the SiPM noise is essential, as it sets the minimum achievable trigger threshold. In particular, correlated noise (DiCT, DeCT, APs) can increase the apparent charge of background signals and raise the accidental trigger rate, limiting how low the threshold can be set. In addition, the SiPM must operate stably at mK temperatures, meeting constraints on power dissipation and thermal load of the dilution fridge, and remaining compatible with the low-noise, low-background environment of the experiment.
@@
In addition to the main dataset used for the characterisation analysis, we also acquired a second, shorter dataset where a small volume of scintillator was coupled to the SiPM. We acquired 2500 waveforms per voltage under the same trigger conditions as before, with a shorter 10 µs acquisition window (1 µs pre-trigger, 9 µs post-trigger). The plastic scintillator used for this study was an early prototype for the SoLid experiment [1], and is reported to have a scintillation efficiency of approximately 10000 photons/MeV e- at room temperature. This was a convenient test sample for the purpose of these initial proof-of-concept measurements.
@@
Three small blocks of the scintillator, each with dimensions of 1 cm (W) $\times$ 3 cm (L) $\times$ 0.5 cm (D), are attached to the inside of a spare copper box lid using epoxy adhesive. The scintillator blocks are each embedded with three wavelength shifting fibres [24], which shift blue light to green with absorption at 420 nm and emission peaking at $\sim$494 nm [1], and guide the photons produced in the scintillator towards the face of the SiPM. Figure 16 shows a picture of the scintillator blocks attached to the inside of the copper box lid. A cross-sectional diagram showing the integration of the scintillator inside the copper box when the lid is attached is shown in Figure 16 (right). The SiPM is not physically in contact with the scintillator blocks when the box is closed; there is an extremely small gap of 1-2 mm between the scintillator blocks and the SiPM face when the box is closed. The SiPM sits in the central area of the three scintillator blocks, for maximum coverage and light collection efficiency.
@@
Figure 16: Left: Image showing the three scintillator blocks attached to the inside of the spare copper box lid. Each block is embedded with three optical fibres, which guide photons generated in the scintillator towards the SiPM surface. Right: Cross-sectional diagram showing integration of scintillator blocks inside copper box when lid is attached and box is closed.
@@
Figure 17 (left) shows an example high-light event observed in this dataset, acquired at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V). In this event, the number of photons incident on the SiPM is so large that it surpasses the dynamic range of the oscilloscope, illustrated by the clipping of the signal at 140 mV. For high photon fluxes such as this, we also observe a lengthening of the signal tail, compared to the SiPM recharge time (<<1 µs). This is caused by the superposition of pulses from real photons and correlated noise, which contribute to an overall increase in the duration of the signal.
@@
Figure 17 (right) compares the pulse area distribution, computed as the 10 µs integral about the event trigger, for data acquired with the scintillator (red) versus data acquired without the scintillator (blue). Without scintillator, we only expect to observe signals associated with the SiPM’s intrinsic noise, resulting in a narrow distribution peaked at the single photon integral. However, once scintillator is introduced, we observe a much broader distribution and a high charge tail, corresponding to energy depositions in the scintillator from particle interactions. This distribution is likely a combination of environmental gamma-rays, which dominate the lower part of the charge distribution, and high-energy cosmic-ray muons populating the high charge tail, as observed in the cryogenic cosmic-ray muon veto system developed by the NUCLEUS experiment [15]. However, this has not been fully characterised here. This data demonstrates that the SiPM is clearly detecting light from the scintillator and optical fibres, which is easily distinguishable above the SiPM noise sources.
@@
As discussed in Section 3.6.1, the SiPM exhibits an increase in noise from long-lasting APs when operated at ultra-low temperatures. Whilst this additional noise source clearly limits the photon counting capability of the SiPM for a low photon flux, this is less of a concern for high photon fluxes generated from cosmic-ray muons. An increase in long-lasting APs will manifest as an increase to the integrated charge of a signal, but only in response to a primary energy deposition in the scintillator. As such, this should not limit the ability of the SiPM to tag cosmic-ray muons through their initial energy deposition. This is a first proof-of-concept of using scintillator coupled to a NUV-HD-cryo SiPM in a mK environment as a means of identifying high-light events consistent with through-going charged particles, such as from cosmic-ray muons.
@@
Figure 17: Left: Example waveform trace taken at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V), illustrating a high-light event consistent with through-going charged particles, potentially from a cosmic-ray muon. Right: Pulse charge distribution (integrated over a 10 µs window) about the event trigger at $V_{\mathrm{bias}}$ = 32 V ($\Delta V$ = 6.3 V) when SiPM is not coupled to scintillator (blue) versus when SiPM is coupled to scintillator (red).
@@ 5 Conclusions
We have shown that a FBK NUV-HD-cryo SiPM is capable of operating successfully inside a cryogen-free dilution refrigerator down to a base temperature of 9.4 $\pm$ 0.2 mK. No measurable increase in the dilution refrigerator base temperature was observed during SiPM operation, indicating that the associated power dissipation is compatible with the available thermal budget. This confirms that the device satisfies the cryogenic cosmic-ray muon veto system requirement for operation directly at the mK stage, coupled to the scintillator, without compromising the fridge performance. We measure the breakdown voltage and gain of the SiPM, and characterise several noise properties including the dark count rate, direct crosstalk probability, and correlated delayed avalanche probability, which we predominately attribute to afterpulsing.
@@
We also present early measurements of deploying the SiPM coupled to scintillator with optical fibre assembly inside the dilution fridge and operated at 9.4 $\pm$ 0.2 mK, as proof-of-concept for a cryogenic cosmic ray muon veto system to be operated internal to a dilution refrigerator. In this run, we observe a significantly broader charge spectrum, consistent with high-light events originating from energy depositions from through-going charged particles, likely from environmental gamma-rays and high-energy cosmic ray muons, which are not present in the data acquired without the scintillator.
The SiPM used in this work originates from an engineering run used for pre-production assembly for the DarkSide-20k experiment and was developed to meet the stringent requirements of single photon detection in a near background-free environment. In contrast, the scintillator-based cryogenic cosmic-ray muon veto system we propose for the QUEST-DMC experiment will operate in a regime of much larger signals, since muons traversing the scintillator will deposit significant energy and produce many photons. The requirements on the noise levels are therefore less stringent than the expected values outlined in [17], provided that muon signals remain sufficiently larger than background contributions from environmental gamma interactions and intrinsic SiPM noise. The system must be able to achieve at least a 90% muon veto efficiency, whilst minimising the experimental dead time due to accidental vetoes.
The dark count rate and direct crosstalk probability we measure in this work appear to be roughly consistent with the expected performance at 77 K for similar $\Delta V$ values [17], which would satisfy the noise requirements for the cosmic-ray muon veto system. These results indicate a weak temperature dependence of these properties between 77 K and 9.4 mK, however this would need to be confirmed with additional measurements at more temperature points across multiple devices. We do however observe a significant increase in the afterpulsing probability, measuring 42% at $\Delta V$ = 6.3 V assuming a 10 µs integration window, compared to an expected afterpulsing probability of 12% at $\Delta V$ = 5 V at 77 K [17]. In addition, we observe evidence of long-lasting afterpulse trains which can last for up to approximately 1 ms, a phenomena which has previously been observed in SiPMs at temperatures below 40 K [18, 27].
While this additional noise does not prevent operation of the veto system, it introduces event-by-event fluctuations in the reconstructed charge and can produce high-charge tails. This increases the probability that background events from environmental gamma interactions exceed the veto threshold, leading to a higher rate of accidental vetoes and hence increased dead time, unless a higher threshold is applied. However, raising the threshold may compromise the ability of the system to achieve the target veto efficiency of 90%. To minimise this excess noise, it will be important to ensure that the scintillation signal is sufficiently prompt, such that the primary light is collected within a short time window. This consideration will inform the next-stage of R&D of the cosmic-ray muon veto system, in particular in the choice of scintillator material, and in optimisation of the geometry and light collection (including the number of required SiPMs and the scintillator-SiPM coupling), to achieve a fast and well-defined muon signal. Following this, we will characterise the cryogenic cosmic-ray muon veto system as a function of temperature, from room temperature down to mK temperatures, in order to determine the minimum veto threshold required to achieve the target performance for the QUEST-DMC experiment.